### v2

# Unleashed from Constrained Optimization: Quantum Computing for Quantum Chemistry Employing Generator Coordinate Inspired Method

Abstract Hybrid quantum-classical approaches offer potential solutions to quantum chemistry problems, yet they also introduce challenges. These challenges include addressing the barren plateau and ensuring the accuracy of the ansätze, which often manifest as constrained optimization problems. In this work, we explore the interconnection between constrained optimization and generalized eigenvalue problems through the Unitary Coupled Cluster (UCC) excitation generators. These generators often serve as building blocks constituting the ansätze in variational quantum eigensolver (VQE) and adaptive derivative-assembled pseudo-Trotter VQE (ADAPT-VQE) simulations. Here, inspired by the generator coordinate method, we employ these UCC excitation generators to construct non-orthogonal, overcomplete many-body generating functions, projecting the system Hamiltonian into a practical working subspace. This approach results in a generalized eigenvalue problem that provides rigorous lower bounds to VQE/ADAPT-VQE energies, effectively bypassing issues related to barren plateaus and heuristic numerical minimizers typical in standard VQE methods. Diverging from conventional quantum subspace expansion methods, we introduce an adaptive scheme that robustly constructs many-body basis sets from a pool of the UCC excitation generators. This scheme supports the development of a hierarchical ADAPT quantum-classical strategy, enabling a balanced interplay between subspace expansion and ansatz optimization to address complex, strongly correlated quantum chemical systems efficiently and cost-effectively. The effective Hamiltonian generated by our approach also supports the computation of excited states and dynamic properties, setting the stage for more advanced quantum simulations in chemistry.
I Introduction
Accurately obtaining ground and excited state energies, along with the corresponding many-body wave functions, is pivotal in comprehending diverse physical phenomena in molecules and materials. This ranges from high-temperature superconductivity in materials like cuprates Imada et al. (1998) and bond-breaking chemical reactions to complex electronic processes in biological and synthetic catalysts with transition metals Witzke et al. (2020) or $f$-block atoms Gould et al. (2022). The associated spin, electronic properties, and dynamics are crucial for deciphering the structure-property-function correlation in various fields, including catalysis, sensors, and quantum materials. However, this task becomes exceptionally challenging in the presence of non-trivial quantum effects, such as strong electron correlation, which influence the evolution of nuclei, electrons, and spins under external stimuli. Conventional wave function methodologies, like configuration interaction, coupled cluster, and many-body perturbation theory, are tailored for diverse electron correlation scenarios Shavitt and Bartlett (2009). Still, they often fall short in handling complex cases or exhibit prohibitive scaling with increasing system size. Consequently, this has become a vigorous area of computational research, encompassing both classical and burgeoning quantum computing studies. The primary objective is to strike an optimal balance between accuracy and computational scalability.
In the realm of quantum computing, significant strides have been made towards promising near-term hybrid quantum-classical strategies, which includes variational quantum algorithm  Peruzzo et al. (2014); McClean et al. (2016); Romero et al. (2018); Shen et al. (2017); Kandala et al. (2017, 2019); Colless et al. (2018); Huggins et al. (2020); Cao et al. (2019); Grimsley et al. (2019a, b); Verteletskyi et al. (2020); McArdle et al. (2019, 2020); Tilly et al. (2021), quantum approximate optimization algorithm Farhi et al. (2014); Stein et al. (2022), quantum annealing Bharti et al. (2021); Albash and Lidar (2018), Gaussian boson sampling Aaronson and Arkhipov (2011), analog quantum simulation Trabesinger (2012); Georgescu et al. (2014), iterative quantum assisted eigensolver McArdle et al. (2019); Motta et al. (2020); Parrish and McMahon (2019); Kyriienko (2020), and many others. These approaches typically delegate certain computational tasks to classical computers, thereby conserving quantum resources in contrast to exclusively to quantum methods. Within this framework, the variational quantum eigensolver (VQE) and its adaptive derivatives are seen as the frontrunners in leveraging near-future quantum advantages Peruzzo et al. (2014); Grimsley et al. (2019a). However, these anticipations are also potentially impeded by the heuristic nature inherent in the critical optimization processes. Issues such as the rigor of ansätze exactness, the challenges in navigating potential energy surfaces replete with numerous local minima, and the numerical nuances in minimization techniques, remain nebulous. These queries are further convoluted when considering the scalability of these methods concerning the number of operators or the depth of quantum circuits involved.
As an alternative to VQE, which employs a highly nonlinear parametrization of the wave function, other near-term strategies aim to construct and traverse a subspace within the Hilbert space to closely approximate the desired state and energy. Typical examples include the quantum subspace expansion Colless et al. (2018); McClean et al. (2017); Takeshita et al. (2020); McClean et al. (2020); Urbanek et al. (2020), the hybrid and quantum Lanczos approaches Suchsland et al. (2021); Motta et al. (2020); Tkachenko et al. (2023), quantum computed moment approaches Seki and Yunoki (2021); Kyriienko (2020); Kowalski and Peng (2020); Vallury et al. (2020); Aulicino et al. (2022), and quantum equation-of-motion approach Ollitrault et al. (2020). These strategies often draw inspiration from truncated configuration interaction approaches or explore the Krylov subspace (see, e.g., Ref. 41 for a recent review on subspace methods for electronic structure simulations on quantum computers). However, the comparison and interrelation of these methods, especially when contrasting subspace expansion with nonlinear optimization, frequently remain unclear. This uncertainty is exacerbated by the choice of ansätze, which rarely guarantees exactness.
Recently, motivated from the generator coordinate approachesHill and Wheeler (1953); Griffin and Wheeler (1957); Rodríguez-Guzmán et al. (2002); Bender et al. (2003); Ring and Schuck (2004); Yao et al. (2010); Egido (2016); Hizawa et al. (2021), we developed a generator coordinate inspired nonorthogonal quantum eigensolver as an alternative near-term approach Zheng et al. (2023), potentially addressing the limitations identified in the VQE/ADAPT-VQE. Similar to its contemporaries, the Generator Coordinate Inspired Method (GCIM) employs low-depth quantum circuits and efficiently utilizes existing ansätze to explore a subspace, targeting specific states and energy levels. Nonetheless, the original GCIM approach demands a priori knowledge of the system to meticulously select both the ansätze and generators, thereby circumventing heuristic approaches but potentially affecting its scalability and efficiency.
The emergence of novel quantum solvers is crucial for facilitating new categories of scalable hybrid algorithms, as exemplified by the recently introduced Quantum Flow (QFlow) algorithm Kowalski and Bauman (2023a). The QFlow algorithm effectively tackles the scalability issue by transforming large-scale problems into coupled eigenproblems that represent smaller active spaces of dimensions. The optimization problems associated with various active spaces can then be efficiently dealt with using considerably fewer quantum resources. Consequently, QFlow approaches are well-suited for developing scalable or distributed quantum algorithms tailored to the sizes of the active spaces and are compatible with existing quantum computers. Scalable QFlow algorithms also have the potential to optimize a large number of wave function parameters simultaneously, due to their distributed nature (see Table 1). To effectively leverage QFlow formulations, active-space solvers must meet specific criteria, namely:
  1. 1.
    cost-effective circuit implementation,
  2. 2.
    availability of efficient solvers providing access to multiple electronic states, and
  3. 3.
    translatability of the GCIM state vector to the unitary coupled cluster (UCC) parameterization.
In the previous paper Zheng et al. (2023), we demonstrated that GCIM algorithms show promise in fulfilling criteria (1) and (2), supplemented by the experiments on an IBM quantum computer presented in Appendix I. This paper will demonstrate that the ground-state vectors estimated by the GCIM framework are the accurate approximation to the exact ground state, and discuss an idea to efficiently translate those approximations into the states prepared by UCC ansätze for the ground-state applications at the end of the paper with two initial results.
Table 1: The total number of unique UCC amplitudes as a function of the size of the target active space, the size of the small-dimensionality active space types included in the QFlow algorithms, and their respective numbers. In columns one and two, we also provide a rough estimate of the number of qubits needed to perform active space optimization. We used $(e_{\rm act},n_{\rm act})$ notation to describe general active space defined by $e_{\rm act}$ active electrons and $n_{\rm act}$ active orbitals.
Target active space QFlow active spaces Number of QFlow Number of optimized
& resources & resources active spaces parameters
(8e,8o) 16 qubits (4e,4o) 8 qubits 36 684
(6e,6o) 12 qubits 16 4,440
(8e,8o) 16 qubits 1 4,899
(10e,10o) 20 qubits (4e,4o) 8 qubits 100 1,775
(6e,6o) 12 qubits 100 19,775
(8e,8o) 16 qubits 25 45,650
(10e,20o) 40 qubits (4e,4o) 8 qubits 1,050 17,325
(6e,6o) 12 qubits 4,550 657,825
(8e,8o) 16 qubits 6,825 7,280,700
(10e,50o) 100 qubits (4e,4o) 8 qubits 9,900 159,975
(6e,6o) 12 qubits 141,900 18,969,975
(8e,8o) 16 qubits 744,975 725,867,100
In this study, we present novel quantum-classical hybrid approaches, inspired by the GCIM approach, aimed at establishing a theoretically exact yet automated subspace expansion procedure. This avoids the optimization limits inherent in conventional VQE/ADAPT-VQE practices. Although integrating straightforward solutions to circumvent these limits—such as designing new bases or applying subspace expansion following constrained optimization—can be achieved Baek et al. (2023); Hirsbrunner et al. (2024); Marti-Dafcik et al. (2024a), domain-specific knowledge in classical quantum chemistry is still required to construct high-quality wave functions. In our approach, by using the conventional Unitary Coupled Cluster (UCC) excitation generators as bases for the subspace, we establish a lower bound to the typically constrained optimization problem in the VQE method. More critically, we introduce a gradient-based, yet optimization-free, automated basis selection method from the conventional UCC operator pool. This enables us to develop a hierarchical ADAPT quantum-classical strategy that allows for an interplay between subspace expansion and ansätz optimization in a controllable manner. Our preliminary results indicate that these new approaches not only excel in addressing challenging, strongly correlated molecular systems but also significantly reducing the simulation time, facilitating deployment on real quantum computers.
II Results and Discussion
We provided a brief review of generator coordinate methods (GCM) in Appendix A. A primary advantage of GCM is that variation occurs in the generating function, rather than directly on the scalar generator coordinate. This distinction enables the variational problem to be addressed through a straightforward eigenvalue process, contrasting with solving a constrained optimization problem, the advantages of which can even be demonstrated with a simple toy model.
Building on this, as depicted in Figure  1, we explore the ground state of a toy model comprising two electrons and four spin orbitals using a VQE approach. The VQE approach is mathematically represented as:
$E_{g}$ $=\min_{\vec{\theta}}\langle\psi_{VQE}(\vec{\theta})|H|\psi_{VQE}(\vec{\theta})\rangle,$ (1)
where $|\psi_{VQE}(\vec{\theta})\rangle$ employs Trotterized UCC single-type ansätze
$|\psi_{VQE}(\vec{\theta})\rangle=G_{2,4}(\theta_{2})G_{1,3}(\theta_{1})|\phi_{0}\rangle$ (2)
with $|\phi_{0}\rangle$ being a reference state and $G_{p_{i},q_{i}}(\theta_{i})=\exp[\theta_{i}(A_{p_{i},q_{i}}-A_{p_{i},q_{i}}^{\dagger})]$ being a single excitation Givens rotation and $A_{p,q}=a_{p}^{\dagger}a_{q}$. The VQE approach here explores a two-dimensional parameter subspace. The consecutive actions of $G_{2,4}$ and $G_{1,3}$ generate four distinct configurations, including the reference. However, due to fewer free parameters than distinct configurations required, the VQE approach is constrained, unable to fully explore the configuration space to guarantee the most optimal solution, irrespective of the numerical minimizer used. In contrast, GCIM solver employs the same Givens rotations, separately or in the product form as discussed by Fukutome Fukutome (1981), to create four generating functions. These functions correspond to a subspace consisting of four non-orthogonal superpositioned states, providing a scope sufficiently expansive to encapsulate the target state. Here, each Givens rotation is essentially a UCC single circuit, simplifying the quantum implementation compared to its double or higher excitation analogs.
From the performance difference we observe that
  • Although the disentangled UCC ansätze can be crafted to be exact Evangelista et al. (2019), the ansätze employed in GCIM approach or in the generalized eigenvalue problem need not be exact. The only prerequisite for the ansätze used in this process is their ability to generate sufficient superpositioned states for a better approximation to the target state to be found in the corresponding subspace.
  • The $N$-qubit Givens rotations ($N\geq 2$) used in both VQE and GCIM solver may originate from the same set, ensuring a comparable level of complexity in the quantum circuits involved. To achieve the same level of exactness, VQE would require a deeper circuit by incorporating additional Givens rotations, while the GCIM algorithm necessitates more measurements (than a single VQE iteration). A direct comparison of the total quantum resources utilized between the two methods would hinge on both the number of generating functions required in the GCIM and the number of iterations needed in VQE.
  • In the context of non-orthogonal quantum eigensolvers, entangled basis sets are usually employed by applying UCC operators on a single determinant reference. Approaches such as the k-UpCCGSD ansätze Lee et al. (2019), UCC-derived ansätze and a cluster-Jastrow modification Baek et al. (2023) have been proposed. These approaches often require domain-specific knowledge available in classical quantum chemistry to construct high-quality wave functions, and rely on the modified Hadamard test of Ref. 12 for evaluating the off-diagonal elements at a polynomial cost. Nevertheless, automating the sophisticated ansatz construction process to facilitate easier deployment of the non-orthogonal quantum solver, especially for non-experts, remains a challenge.
These observations then lead us to the pivotal question in advancing a more efficient GCIM algorithm for both classical and quantum computation: “How can we efficiently select the generating functions to construct a subspace in the Hilbert space that encompasses the target state?
Figure 1: (a) A toy model consists of two-electron (one alpha electron and one beta electron) in four spin-orbitals (two alpha spin-orbitals and two beta spin-orbitals), where the spin-flip transition is assumed forbidden. (b) The projection of the exact wave function on each configuration (yellow shadow). A constrained optimization of the free parameters will put limits on the projection (dashed line). (c) Comparative demonstration between GCIM method and VQE. The GCIM method generates a set of non-orthogonal bases, called generating functions. Then, the GCIM explores the projection of the system on these generating functions and solves a corresponding generalized eigenvalue problem for the target state and its energy. The conventional VQE essentially explores the parameter subspace for a given wave function ansatz through numerical optimization that can be usually constrained by many factors ranging from ansatz inexactness to barren plateaus and others. For given Givens rotations that generated excited state configurations, the lowest eigenvalue obtained from the GCIM method guarantees a lower bound of the most optimal solution from the VQE. It is worth mentioning that: (i) for a standard (single-circuit) VQE, having one two-qubit Givens rotation acting on a standard quantum-chemistry reference (restricted Hartree-Fock) would not improve the energy estimate; (ii) “fermionic swap” gatesKivlichan et al. (2018) would be required if the excitations included in the ansätze involve spin-orbitals that are not mapped onto adjacent qubits for a given quantum architecture.
Figure 2: Number of generating functions as functions of the number of Givens rotations (denoted as $m$) included in the GCIM wave function ansätze at different truncation levels.
In constructing a non-trivial UCC ansatz for a molecular system comprising $n_{e}$ electrons in $N$ spin orbitals, one might consider applying a sequence of $K\leq n_{e}$ Givens rotations to a reference state $|\phi_{0}\rangle$ to generate all the possible configurations. For example, if each Givens rotation corresponds to a single excitation, the disentangled UCC Singles ansatz can be expressed as:
$|\psi(\vec{\theta})\rangle=\prod_{i=1}^{K}G_{p_{i},q_{i}}(\theta_{i})|\phi_{0}\rangle,$ (3)
where each $G_{p_{i},q_{i}}$ generates a superposition of no more than two states. Therefore, the total number of configurations, $n_{c}$, within the superpositioned ansätze (3)--where $\vec{\theta}=\{\theta_{i}|i=1,\cdots,K\}$ varies--cannot exceed $2^{K}$. This number indicates the maximum number of generating functions utilized to achieve the most optimal solution within the corresponding subspace but exhibits an exponential scaling relative to the count of Givens rotations applied. A strategy to potentially generate the full set of $2^{K}$ generating functions involves applying $1,2,\cdots,N$ two-qubit Givens rotations separately to the reference, as shown below:
  • $|\phi_{0}\rangle,$
[…]
  • $\prod_{i=1}^{K}G_{p_{i},q_{i}}(\theta_{i})|\phi_{0}\rangle$
Here, the number of generating functions employing $k$ Givens rotations is equivalent to the number of $k$-combinations from the set of $K$ Givens rotations, aligning with the combinatorial identity $\sum_{k=0}^{K}\bigg{(}\begin{array}[]{c}K\\ k\end{array}\bigg{)}=2^{K}$. Importantly, when $K=n_{e}$, each sequence of Givens rotations in the above list can be tailored to include a unique excited configuration within the associated subspace, thereby mapping the entire configuration space to ensure exactness in this limit. This characteristic bears notable similarity to full configuration interaction treatments, with the distinctive aspect that this approach constructs a non-orthogonal many-body basis set. In practice, to capture strong correlation with a fewer number of generating functions, generalized $N$-qubit Givens rotations are commonly applied in quantum chemical simulations. A typical example includes the particle-hole one- and two-body unitary operators frequently employed in VQE and ADAPT-VQE simulations. Notably, it has been demonstrated that products of only these particle-hole one- and two-body unitary operators can approximate any state with arbitrary precision.Evangelista et al. (2019)
It’s important to recognize that the generating functions, including the ansätze, are not generally orthogonal. This non-orthogonality necessitates extra resources for evaluating the associated off-diagonal matrix elements and the entire overlap matrix for solving a generalized eigenvalue problem. Additionally, if higher-order Givens rotations are involved, the configuration subspace expansion will be influenced by the ordering of the Givens rotations in the ansätze. Nevertheless, the ordering limitation can be partially circumvented by allowing the possible permutations of the Givens rotations within the generating functions.
The strategy outlined above can also be employed to establish a hierarchy of approximations to the GCIM wave function, targeting a configuration subspace where the ansätz (3) is fully projected out. For example, the approximation at level $k$ ($0\leq k<K$) can be defined by a working subspace that incorporates both the ansätz (3) and some generating functions produced by acting a product of at most $k$ Givens rotations on the reference. Notably, when $k=K$, regardless of the rotations in the generating functions, the configuration subspace has expanded sufficiently for the ansätz (3) to be fully projected out, completely removing the constraints in the optimization. For other approximations where $0\leq k<K$, a variational procedure can be undertaken before the configurational subspace expansion, aiming to deduce a lower bound for the optimal expectation value of $H$ with respect to the ansätz (3). In this fashion, the GCIM solver can be implemented as a subsequent step either after each VQE iteration or following a complete VQE calculation to unleash some constraints due to the local minima and the inexactness of the ansätz.
Figure 3: (a) Schematic depiction of the ADAPT-GCIM algorithm in $(n+1)$-th iteration. The working subspace and its corresponding configuration subspace are characterized by a surrogate state $|\psi\rangle_{s}$. (b) In a two-configuration subspace, a Givens rotation can couple two configurations, effectively generating a surrogate state, a superposition of two configurations, that can be employed as a metric to characterize this subspace. The quality of the working subspace, comprising one configuration and the surrogate state, is not sensitive to the specific choice of scalar rotation.
To facilitate a more robust selection of generating functions, we consider the fluctuation of the GCIM energy when a new generating function is included. In the context of VQE, similar considerations have led to the development of its ADAPT (Adaptive Derivative-Assembled Pseudo-Trotter) version. In the GCIM framework, several routines can be proposed for a gradient-based ADAPT-GCIM approach. A straightforward method is to directly compute the GCIM energy gradient with respect to the scalar rotation, as detailed in Appendix C. The energy gradient computed in this way depends on the eigenvector solved from the GCIM general eigenvalue problem after the inclusion of new generating functions. However, when solving a generalized eigenvalue problem, the eigenvalue (i.e. the energy) is more sensitive to the working subspace than to the scalar rotation in the employed generating functions. This sensitivity is exemplified in the Lemma 1 of Theorem 1 in Appendix B, where, for a $2\times 2$ generalized eigenvalue problem, the choice of the scalar rotations in the employed generating functions is relatively flexible.
This consideration leads us to propose a more flexible, gradient-based ADAPT-GCIM algorithm (illustrated in Figure 3a). In this approach, we use a surrogate product state $|\psi\rangle_{s}$ that, irrespective of its specific non-singular scalar rotations, primarily serves as a metric to (i) abstract the change in the GCIM working subspace and (ii) approximate the GCIM energy gradient calculation. Utilizing this metric, we can compute the gradient in a manner similar to that in the conventional ADAPT-VQE. However, a significant difference from ADAPT-VQE is that in ADAPT-GCIM, we skip the optimization of phases, which as we will see later significantly reduces the simulation time. Also, the energy in ADAPT-GCIM is not directly obtained from the measurement of the expectation value with respect to the surrogate state. Instead, it is obtained from solving a generalized eigenvalue problem within the working subspace. Additionally, it is worth noting that both optimization and solving the eigenvalue problem can be integrated into our proposed ADAPT-GCIM framework to explore a better balance between computational cost and efficiency. In this regard, we propose the ADAPT-GCIM($m,n$), detailed in Appendix H, where the optimization is activated for at most $n$ rounds in every $m$ ADAPT-GCIM iterations. As we will see later, in comparison with the ADAPT-GCIM, the ADAPT-GCIM($m,n$) approach allows for adjustment of the $m$ and $n$ values to reduce the number of iterations while slightly increasing the computational cost and simulation time.
Figure 4: Convergence performance of ADAPT-GCIM in computing the ground states of seven molecules. For comparison, results from ADAPT-VQE, ADAPT-VQE-GCIM, and ADAPT-VQE-GCIM(1) are also included. The shaded region is when error is smaller than the chemical accuracy. ADAPT-VQE and ADAPT-VQE-GCIM are terminated when the sum of the magnitudes of all gradients falls below $10^{-4}$. ADAPT-GCIM is set to converge if the energy differences are under $10^{-6}$ hartrees for a certain number of consecutive iterations, where this number depends on the size of the operator pool.
We have performed numerical experiments on four molecular systems across seven geometries to test the performance of our proposed ADAPT-GCIM approach in searching the ground states corresponding to different electronic configurations. For instance, the more pronounced quasidegeneracy in the almost-square $H_{4}$ compared to linear $H_{4}$ Zheng et al. (2023), and in stretched $H_{6}$ as opposed to compact $H_{6}$, presents strong static correlations that pose a challenge for traditional single-reference methods Jankowski and Paldus (1980). Additionally we proposed two GCIM modified ADAPT-VQE approaches: ADAPT-VQE-GCIM and ADAPT-VQE-GCIM(1) for performance analysis. In the ADAPT-VQE-GCIM approach, a GCIM step is performed after each ADAPT-VQE iteration, while in the ADAPT-VQE-GCIM(1) approach, a ‘one-shot’ GCIM calculation is performed at the conclusion of an ADAPT-VQE calculation (see Appendix D for more details). The performance results, as shown in Figure 4, include comparisons with ADAPT-VQE. It is evident that the GCIMs offer several strategies for locating a lower bound to the ADAPT-VQE energy. In all the test cases, the ADAPT-VQE-GCIM and ADAPT-GCIM approaches exhibit steady monotonic converging curves, almost always reaching close to machine precision. The ADAPT-VQE-GCIM(1) also slightly improves the ADAPT-VQE energy upon convergence. Notably, the ADAPT-VQE-GCIM consistently achieves a lower energy than ADAPT-VQE at every single iteration, thereby converging faster than ADAPT-GCIM. As a trade-off for skipping the phase optimization, the ADAPT-GCIM typically requires more iterations to converge compared to ADAPT-VQE and ADAPT-VQE-GCIM. However, both ADAPT-GCIM and ADAPT-VQE-GCIM are capable of achieving numerically exact solutions for almost all the testing cases.
The ADAPT-GCIM approximation to the ground-state vector of $H$ is reconstructed by
|$\ket{\psi_{GCIM}}=\mathcal{N}\cdot\sum_{j}f_{j}\ket{\psi_{j}},$
where $f$ is the eigenvector solved from the generalized eigenvalue problem in ADAPT-GCIM at the final iteration and {|$\{\ket{\psi_{j}}\}$ is the selected basis set. $\mathcal{N}$denotes the normalization factor. Table 2 provides the overlaps between the exact ground-state vector, |$\ket{\psi_{exact}}$, and corresponding GCIM approximation for all molecules in this work, numerically proving the accuracy of the ADAPT-GCIM method.
Table 2: Overlaps between ADAPT-GCIM and exact ground-state vectors, machine epsilon is $2.22\times 10^{-16}$.
Molecule $1-|\braket{\psi_{GCIM}}{\psi_{exact}}|^{2}$
$H_{4}$ (linear) $2.22\times 10^{-16}$
$H_{4}$ (square) $<2.22\times 10^{-16}$
$LiH$ $2.22\times 10^{-16}$
$BeH_{2}$ $2.22\times 10^{-15}$
[…]
Table 3: Excitation energies in eV from different methods
Molecule Character of EOMCC- EOMCC- ADAPT-
excited state SD SDT GCIM
$H_{4}$ (linear) SD(1 excited) 12.738 12.555 12.565
$H_{4}$ (square) D (1 excited) 4.275 4.273 4.183
$LiH$ SD (1 excited) 3.588 3.586 3.586
$BeH_{2}$ SD (3 excited) 15.558 14.563 14.545
[…]
In assessing the quantum resources needed for the ADAPT approaches, ADAPT-VQE-GCIM and ADAPT-VQE-GCIM(1) require additional resources compared to ADAPT-VQE approach to improve energy results. The number of CNOT gates for ADAPT-VQE and ADAPT-GCIM depends on the preparation of the operators and the number of iterations to reach desired accuracy. Figure 5 demonstrates the total number of CNOTs for the newly formed ansätze at each energy error level. To avoid the confusion, the term “newly formed ansätze” is defined as the following:
  • for ADAPT-VQE, the newly formed ansätz is the matrix product $\prod_{i=1}^{k}G^{VQE}_{i}(\theta_{i})|\phi_{\rm HF}\rangle$ at iteration $k$ that prepares the VQE state |$\ket{\psi_{VQE}}$,
  • for ADAPT-GCIM, the newly formed ansätze are $G^{GCIM}_{k}(\theta_{k})|\phi_{\rm HF}\rangle$ and $\prod_{i=1}^{k}G^{GCIM}_{i}(\theta_{i})|\phi_{\rm HF}\rangle$ at iteration $k$,
where $G^{VQE}_{i}$ and $G^{GCIM}_{i}$ are UCC single or double excitation operators.
Figure 5: Energy difference as a function of the number of CNOT gates in the circuits of newly formed ansätze. For ADAPT-VQE, such ansatz is the product of operators that prepare the VQE state |$\ket{\psi_{VQE}}$. For ADAPT-GCIM, the newly formed ansätze are the operators the newly selected Givens rotation operator and the product of Givens rotation operators. All circuits are generated from the order-1 trottered fermionic operators by Qiskit Qiskit contributors (2023). The reduced counts are obtained using qiskit.transpile with optimization level 3, which includes canceling back-to-back CNOT gates, commutative cancellation, and unitary synthesis. Our estimations to ADAPT-VQE match with the corresponding calculations in Ref. 63. A further discussion on the calculation of the number of CNOT gates is in Appendix E.
[…]
Table 4: Simulation costs for energy calculations of seven geometries using the ADAPT approaches on a laptop equipped with an Apple M3 Max chip. During the optimization rounds, the default optimizer is the Broyden–Fletcher–Goldfarb–Shanno (BFGS) algorithm, implemented in SciPy Virtanen et al. (2020). ADAPT-GCIM(5,2) is an extended method introduced in Appendix H.
Molecule ADAPT- # of ADAPT Total # of Optim- Total Simulation Time (s) Ground-State
Iterations ization Rounds Gradients Energy Eval. Energy Error (in a.u.)
$H_{4}$ (linear) GCIM 24 0 0.16 2.05 7.55 ×$\times 10^{-15}$
VQE 11 121 0.10 7.61 7.11 ×$\times 10^{-15}$
$H_{4}$ (square) GCIM 26 0 0.17 2.49 2.22 ×$\times 10^{-15}$
VQE 11 136 0.09 17.44 4.88 ×$\times 10^{-15}$
$LiH$ GCIM 81 0 3.67 75.09 1.95 ×$\times 10^{-14}$
VQE 27 439 1.24 253.64 6.36 ×$\times 10^{-8}$
$BeH_{2}$ GCIM 112 0 10.78 283.67 1.07 ×$\times 10^{-14}$
VQE 41 1248 3.69 2308.14 1.91 ×$\times 10^{-06}$
[…]
VQE 70 4792 3.17 4627.37 1.70 ×$\times 10^{-6}$
GCIM(5,2) 89 36 4.36 162.21 6.66 ×$\times 10^{-15}$
[…]
VQE 88 9209 3.99 11924.84 9.98 ×$\times 10^{-08}$
[…]
VQE 84 11313 3.86 12765.16 1.15 ×$\times 10^{-6}$
GCIM(5,2) 51 22 2.69 47.76 9.61 ×$\times 10^{-8}$
Regarding measurements in fault tolerant quantum computation, ADAPT-VQE requires measurements for VQE and gradient calculations at each iteration. The number of VQE measurements depends on the average optimizations per iteration ($\tilde{N}_{\rm opt}$) and the total number of iterations ($N_{\rm iter}$). Gradient measurements are roughly equal to the size of the pool multiplied by the number of unitary terms in the Hamiltonian ($N_{\rm term}$), which can be reduced with grouping techniques. Thus, total measurements for ADAPT-VQE scales as $\mathcal{O}(\tilde{N}_{\rm opt}\times N_{\rm iter}+N_{\rm term}\times N_{\rm iter})$. In ADAPT-GCIM, measurements are needed for gradients and constructing the $\mathbf{H}$ and $\mathbf{S}$ matrices. The latter scales as the squared number of generating functions, $\mathcal{O}(N_{\rm GF}^{2})$, but per iteration, it only grows as $\mathcal{O}(N_{\rm GF})$, assuming a linear increase of generating functions with iterations. Since $N_{\rm GF}\sim\mathcal{O}(N_{\rm iter})$, measurement comparison between the two ADAPT approaches reduces to $\mathcal{O}(\tilde{N}_{\rm opt}\times N_{\rm iter(VQE)})$ vs. $\mathcal{O}(N_{\rm iter(GCIM)}^{2})$.
In practice, real quantum measurements often encounter significant noise, especially in solving the generalized eigenvalue problem where finite shot uncertainty is a concern Zheng et al. (2023). Stronger correlations typically require substantially more measurements than weaker ones. As detailed in Appendix G, errors in the $\mathbf{S}$ matrix have a greater impact than those in the $\mathbf{H}$ matrix. However, with importance sampling, accuracy can still be improved by orders of magnitudes in most cases despite these challenges (see Figure 7). We deployed the ADAPT-GCIM computation of the strongly correlated linear H4 molecule on an IBM superconducting quantum computer, $ibm\_osaka$, to conduct a rudimentary test of the applicability of the GCIM approaches on emerging hardware. As shown in Appendix I, with problem-specific error mitigation methods, the error in ground-state energy estimation was reduced from 0.046 a.u. to $3.9\times 10^{-9}$ a.u.
III Conclusion and Outlook
In this study, we have introduced a novel hybrid quantum chemical approach that enhances the accuracy and efficiency of quantum simulations. This method diverges from typical constrained optimization methods like VQE and its ADAPT variant, representing a new subspace expansion method that employs non-orthogonal Givens rotations as the basis for expansion. Our approach automates basis selection, in contrast to other subspace expansion methods. Unlike the ADAPT-VQE, our method, while using the same ansätze pool, solves a generalized eigenvalue problem within a working subspace, where the global minimum provides a lower bound to the ADAPT-VQE solutions. Numerically, the ADAPT-GCIM has shown superior performance over the ADAPT-VQE, especially in cases of strong electron correlation.
The GCIM approach stands out among other subspace expansion and generalized eigenvalue problem-solving methods due to its size-extensiveness, attributed to its use of UCC ansätze as bases. The approach’s scalability for larger systems will be enhanced by its direct application to qubit space, using the exponential of anti-hermitian Pauli strings as generating functions. This strategy aligns with developments like qubit-ADAPT-VQE, which offers scalability for the ADAPT-VQE method Tang et al. (2021).
The straightforwardness of the GCIM ansätze makes it a promising solver also for excited state computation and Hamiltonian downfolding, as demonstrated in the ADAPT-GCIM procedure (Figure 3a). This lays the foundation for advanced quantum simulations in chemistry. Furthermore, its potential in designing constant-circuit-depth quantum algorithms, such as the QFlow algorithm Kowalski and Bauman (2023b), is significant. However, integrating the GCIM solver into QFlow will require robust methods to approximate GCIM state, |$\ket{\psi_{GCIM}}$, by UCC-type ansätze. That is, we need to find |$\ket{\psi_{UCC}^{*}}:=\ket{\psi_{UCC}(\vec{\theta}^{*})}$ where
$\vec{\theta}^{*}:=\operatorname*{arg\,max}_{\vec{\theta}}\,\,|\braket{\psi_{GCIM}}{\psi_{UCC}(\vec{\theta})}|^{2}.$
The most straightforward scheme is to use the UCCSD ansatz. We illustrate the feasibility through seven examples in Table 5, where $E_{exact}$ is the exact ground-state energy and $E_{UCCSD}$ is its approximation from |$\ket{\psi_{UCCSD}^{*}}$. For future implementation, a possible direction is to employ an overlap-based adaptive algorithm as described in Ref. 66 to improve the efficiency and robustness.
Table 5: Overlap between GCIM state and the optimized UCCSD states, and the ground-state energy error of the UCCSD states.
Molecule $1-|\braket{\psi_{GCIM}}{\psi_{UCCSD}^{*}}|^{2}$ $|E_{UCCSD}-E_{exact}|$
(in a.u.)
[…]
$LiH$ $3.76\times 10^{-6}$ $1.10\times 10^{-5}$
$BeH_{2}$ $2.89\times 10^{-4}$ $4.06\times 10^{-4}$
[…]
Last but not least, while the gate counts in Figure 5 match with the estimations for fermionic operators in Ref. 63, the CNOT counts can be greatly reduced by the implementations in other works in the future Berry et al. (2007); Anselmetti et al. (2021); Arrazola et al. (2022); Kottmann (2023). Another important direction to explore is to evade Hadamard test for evaluating off-diagonal matrix entries in ADAPT-GCIM to effectively reduce the hardware requirements in the circuit implementationHuggins et al. (2020); Cortes and Gray (2022); Motta et al. (2024b).
IV Method and Computational Details
The operator pool employed in all the ADAPT simulations in the present study consists of generalized singly and doubly spin-adapted excitation operators. In all the ADAPT-VQE simulations, the convergence is achieved when the norm of the gradient vector is less than $10^{-4}$. Here, we use the more strict criteria than in Ref. 73 to show the full picture of the performance of ADAPT-VQE. Other criteria, such as the magnitude of the change of the lowest eigenvalue, can be used for earlier convergence with the assistance of ADAPT-VQE-GCIM. More details are listed in Appendix F.
One issue in the GCIM simulations is the numerical instability of solving the generalized eigenvalue problem. It is possible that the overlap matrix $\mathbf{S}$ becomes numerically indefinite at certain iterations. In classical computing, the issue of singularity can be perfectly avoided by orthogonalizing the generating function basis set. However, in quantum computing, the quantum version of the orthogonalization algorithm like the quantum Gram-Schmidt process usually requires extra quantum resources, such as QRAM Zhang et al. (2021). In the present study, when a singular overlap matrix is encountered, we choose to project both $\mathbf{H}$ and $\mathbf{S}$ using the eigenvectors of $\mathbf{S}$ corresponding to large positive eigenvalues Urbanek et al. (2020). Without the disturbance due to the finite number of shots, the threshold for a sufficiently large eigenvalue can be as relaxed as $10^{-13}$. Otherwise, as introduced in Appendix G, it is usually set to $10^{-5}$ and $10^{-6}$. A comparison test in Appendix F shows that this mitigation is very effective in achieving the accuracy levels reachable by GCIM methods.
In the proposed ADAPT-VQE-GCIM(1), the number of bases in the working subspace equals to the number of Givens rotations in the ansatz generated at the last ADAPT-VQE iteration, incremented by one. This increment accounts for the product of these Givens rotations corresponding to the ansatz itself. Conversely, in the ADAPT-GCIM and ADAPT-VQE-GCIM approaches, the number of bases is twice the number of iterations. To illustrate this, consider the $k$-th iteration: when the Givens rotation $G_{k}$ with the greatest gradient is selected, two associated basis vectors,
  • $G_{k}(\theta_{k})|\phi_{\rm HF}\rangle$ and
  • $G_{k}(\theta_{k})\prod_{i=1}^{k-1}G_{i}(\theta_{i})|\phi_{\rm HF}\rangle$
are added to the working subspace. So the number of bases is increased by two in every iteration. The key difference is that in ADAPT-GCIM, the phase is set to a constant scalar (e.g., $\pi/4$) for any Givens rotation in the ansätz, while in ADAPT-VQE-GCIM, the phases $\theta_{1}$ to $\theta_{k}$ are optimized during the $k$-th iteration. In all the experiments of ADAPT-GCIM, $\theta_{k}$ is set to $\pi/4$ for all iterations in seven geometries.
The implementation of ADAPT-VQE is the original code in Ref. 73 from the corresponding repository Grimsley et al. (2022), and its computation is based on the SciPy sparse linear algebra package Virtanen et al. (2020). To have the same foundation for comparison, calculations in Figure 4 for three GCIM algorithms follow the same manner. The quantum resources estimation, shown in Figure 5, utilizes Qiskit Qiskit contributors (2023) for trotterization and circuit generation for excitation operators. All the numerical experiments were carried on a laptop with an Apple M3 Max chip.
V Author Contributions
B. P. conceptualized the idea and designed the research, M. Z. and B. P. implemented the approaches and performed the simulations. All authors contributed to paper writing.
VI Acknowledgements
M. Z., B. P. and K. K. were supported by the “Embedding QC into Many-body Frameworks for Strongly Correlated Molecular and Materials Systems” project, which is funded by the U. S. Department of Energy, Office of Science, Office of Basic Energy Sciences (BES), the Division of Chemical Sciences, Geosciences, and Biosciences under FWP 72689. A. L. was supported by the U.S. Department of Energy, Office of Science, National Quantum Information Science Research Centers, Quantum Science Center (QSC). K. K. also acknowledges the support from Quantum Science Center (QSC), a National Quantum Information Science Research Center of the U.S. Department of Energy (under FWP 76213). X. Y. was supported by National Science Foundation CAREER DMS-2143915. M. Z. and X. Y. both also were supported by Defense Advanced Research Projects Agency as part of the project W911NF2010022: The Quantum Computing Revolution and Optimization: Challenges and Opportunities. This research used resources of the Oak Ridge Leadership Computing Facility, which is a DOE Office of Science User Facility supported under Contract DE-AC05-00OR22725. The Pacific Northwest National Laboratory is operated by Battelle for the U.S. Department of Energy under Contract DE-AC05-76RL01830.
VII Competing interests
All authors declare no financial or non-financial competing interests.
VIII Data and code availability
The data and the code for the related numerical simulation is available in the GitHub repository Zheng et al. (2022).
Appendix A Generator Coordinate: A Brief Overview
Originating from studies of nuclear structure and nuclear reactions, including those occurring in heavy-ion collisions or astrophysical scenarios, the Generator Coordinate Method (GCM), proposed by Griffin-Hill-Wheeler (GHW) Hill and Wheeler (1953); Griffin and Wheeler (1957), approximates the many-body wave function $|\Psi\rangle$ of the Hamiltonian $H$ by describing collective motion through the variation of one or more collective generator coordinates, denoted as ${\vec{\alpha}}=\{\alpha_{i}\}$, where $i=1,\cdots,N_{\alpha}$ and $N_{\alpha}>1$. This variation leads to the representation of the many-body wave function of a quantum system in terms of a set of basis states (also called generating functions) $\{|\phi(\vec{\alpha})\rangle\}$:
$|\Psi\rangle\approx\int{\rm d}{\vec{\alpha}}|\phi({\vec{\alpha}})\rangle f({\vec{\alpha}}).$ (4)
Here, $f({\vec{\alpha}})$ is an unknown ${\vec{\alpha}}$-dependent weight function. The choice of these generator coordinates ${\vec{\alpha}}$ depends on the specific system under study. The fundamental concept of the GCM is to treat these generator coordinates as variational parameters, which are then used to approximate the target wave function as a superposition of $\{|\phi({\vec{\alpha}})\rangle\}$. The GCM’s ability to adapt to changing collective coordinates and its variational nature make it well-suited for describing complex states that break symmetries, allowing for a flexible and efficient description of properties such as ground states, excited states, and transitions between them.
Technically, to determine $f({\vec{\alpha}})$, substituting (4) into the Schrödinger equation, and then the variational process leads to the so-called GHW integral equation:
$\int{\rm d}{\vec{\alpha}}^{\prime}\bigg{(}\langle\phi(\vec{\alpha})|H|\phi(\vec{\alpha}^{\prime})\rangle-E\langle\phi(\vec{\alpha})|\phi(\vec{\alpha}^{\prime})\rangle\bigg{)}f({\vec{\alpha}}^{\prime})=0.$ (5)
This integral equation can be discretized in a finite-dimensional subspace of the many-body Hilbert space, leading to a general eigenvalue problem:
$\sum_{j=1}^{N_{\alpha}}\bigg{(}\mathbf{H}_{ij}-E\mathbf{S}_{ij}\bigg{)}f(\alpha_{j})=0,$ (6)
where
$\mathbf{H}_{ij}$ $=\langle\phi(\alpha_{i})|H|\phi(\alpha_{j})\rangle,$
$\mathbf{S}_{ij}$ $=\langle\phi(\alpha_{i})|\phi(\alpha_{j})\rangle.$
While the GCM is a powerful method, it can be computationally demanding, especially when dealing with a large number of basis states or when symmetry-breaking effects are intricate. For strongly correlated molecular systems, these requirements are often associated with incorporating sufficient electron correlation effects into GCM calculations, which typically involves using more sophisticated electronic structure methods than mean-field theory. Techniques such as configuration interaction (CI), coupled-cluster (CC) methods, and many-body perturbation theory (MBPT) can be employed to capture electron correlation effects accurately Shavitt and Bartlett (2009). The basis functions generated from these methods should be capable of representing the electron correlation effects relevant to the quantum phenomena under investigation. Various levels of electron correlation can be included in GCM calculations, depending on the chosen basis functions and computational resources available. It is interesting to observe the GCM as a generalization of the non-orthogonal determinants (e.g., non-orthogonal configuration interaction). Even though non-entangled single-particle states were employed in the GCM for simplifying the implementation in our previous report Zheng et al. (2023), the entangled many-body states can also be employed in the framework to involve complex correlations among particles, making it suitable for describing systems with complex shapes and collective motion.
Appendix B Mapping optimization to a configuration subspace
Theorem 1. For a Hamiltonian $H$ and a parametrized normalized ansatz $|\Phi(\vec{\theta})\rangle=U(\vec{\theta})|\phi\rangle$, where $|\phi\rangle$ is a fixed reference state and $U(\vec{\theta})$ is a unitary operator dependent on the parameter $\vec{\theta}=(\theta_{1},\theta_{2},\cdots\theta_{P})$, the minimal expectation value
$\epsilon=\min_{\vec{\theta}}\langle\phi|U^{\dagger}(\vec{\theta})|H|U(\vec{\theta})|\phi\rangle$ (7)
is the upper bound of the minimal eigenvalue of $H$ in an arbitrary linearly independent $M$-configuration subspace $\{\psi_{1},\psi_{2},\cdots,\psi_{M}\}$ ($M\geq P$) where $U(\vec{\theta})|\phi\rangle$ can be completely projected out, i.e.,
$U(\vec{\theta})|\phi\rangle=\sum_{i=1}^{M}c_{i}(\vec{\theta})|\psi_{i}\rangle.,$ (8)
where
$c_{i}(\vec{\theta})=\sum_{j=1}^{M}(\mathbf{S}^{-1})_{ij}\langle\psi_{j}|U(\vec{\theta})|\phi\rangle$ (9)
and $\mathbf{S}$ is the the overlap matrix with
$\mathbf{S}_{ij}=\langle\psi_{i}|\psi_{j}\rangle.$ (10)
Proof : Given (8), $U(\vec{\theta})|\phi\rangle$ is one state in the linearly independent $M$-configuration subspace, thus the minmal expectation value from minimizing (7) is constrained by the form of $U(\vec{\theta})$, and does not necessarily give the minimal expectation value of $H$ in the subspace. Instead, searching the minimal expectation value of $H$ in the subspace can be re-formulated to minimizing a generalized Rayleigh quotient for the projected Hamiltonian matrix $\mathbf{H}$ and the overlap matrix $\mathbf{S}$
$\min_{f}\leavevmode\nobreak\ R(f)=\frac{f^{\ast}\mathbf{H}f}{f^{\ast}\mathbf{S}f},\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \text{subject to}\leavevmode\nobreak\ \leavevmode\nobreak\ f^{\ast}\mathbf{S}f=1,$ (11)
with matrix elements defined as
$\mathbf{H}_{ij}$ $=\langle\psi_{i}|H|\psi_{j}\rangle,$ (12)
$\mathbf{S}_{ij}$ $=\langle\psi_{i}|\psi_{j}\rangle.$ (13)
By introducing a Lagrange multiplier and forming the Lagrangian,
$L(f,\epsilon)=f^{\ast}\mathbf{H}f-\epsilon(f^{\ast}\mathbf{S}f-1),$ (14)
the minimization corresponds to taking the derivative of $L(f,\epsilon)$ with respect to $f$ and setting it to zero, leading to a generalized eigenvalue equation:
$\mathbf{H}f=\epsilon\mathbf{S}f,$ (15)
where the lowest eigenvalue corresponds to the global minimum in the $M$-configuration subspace.
Lemma 1 Consider a unitary operation $U$ parametrized by a single scalar $\theta$ in a variational ansätze. If, for any value of $\theta$, the action of $U$ on $|\phi\rangle$ produces at most two distinct configurations, then the task of minimizing $\langle\phi|U^{\dagger}(\vec{\theta})|H|U(\vec{\theta})|\phi\rangle$ can be equivalently approached by solving the following $2\times 2$ generalized eigenvalue problem:
$\left(\begin{array}[]{cc}\langle\phi|U^{\dagger}(\theta_{1})HU(\theta_{1})|\phi\rangle&\langle\phi|U^{\dagger}(\theta_{1})HU(\theta_{2})|\phi\rangle\\ \langle\phi|U^{\dagger}(\theta_{2})HU(\theta_{1})|\phi\rangle&\langle\phi|U^{\dagger}(\theta_{2})HU(\theta_{2})|\phi\rangle\end{array}\right)f$ (18)
$\leavevmode\nobreak\ \leavevmode\nobreak\ =\epsilon\left(\begin{array}[]{cc}1&\langle\phi|U^{\dagger}(\theta_{1})U(\theta_{2})|\phi\rangle\\ \langle\phi|U^{\dagger}(\theta_{2})U(\theta_{1})|\phi\rangle&1\end{array}\right)f.$ (21)
Here $\theta_{1}$ and $\theta_{2}$ are chosen such that $\|\langle\phi|U^{\dagger}(\theta_{1})U(\theta_{2})|\phi\rangle\|^{2}\neq 1$. The lowest eigenvalue of the $2\times 2$ problem yields the minimal energy within the defined subspace.
Proof : For any $\theta$, let’s express
$U(\theta)|\phi\rangle=a(\theta)|\phi_{1}\rangle+b(\theta)|\phi_{2}\rangle$ (22)
where $|\phi_{1}\rangle$ and $|\phi_{2}\rangle$ are orthogonal, i.e., $\langle\phi_{1}|\phi_{2}\rangle=0$, and $a(\theta)$ and $b(\theta)$ are scalar functions of $\theta$ with $|a(\theta)|^{2}+|b(\theta)|^{2}=1$
According to Theorem 1, the minimization of $\langle\phi|U^{\dagger}(\vec{\theta})|H|U(\vec{\theta})|\phi\rangle$ over all $\theta$ can be recast into a $2\times 2$ eigenvalue problem with the Hamiltonian matrix
$\mathbf{H}=\left(\begin{array}[]{cc}\langle\phi_{1}|H|\phi_{1}\rangle&\langle\phi_{1}|H|\phi_{2}\rangle\\ \langle\phi_{2}|H|\phi_{1}\rangle&\langle\phi_{2}|H|\phi_{2}\rangle\end{array}\right).$ (25)
Alternatively, the expansion of $U(\theta)|\phi\rangle$ can also be expanded by non-orthogonal basis, e.g.
$U(\vec{\theta})|\phi\rangle=c_{1}U(\theta_{1})|\phi\rangle+c_{2}U(\theta_{2})|\phi\rangle,$ (26)
where $\theta_{1}$ and $\theta_{2}$ are two scalars satisfying $\|\langle\phi|U^{\dagger}(\theta_{1})U(\theta_{2})|\phi\rangle\|^{2}\neq 1$, and $c_{1}$ and $c_{2}$ depends on the choice of $\theta,\theta_{1},\theta_{2}$ and can be obtained through
$(a(\theta_{1})b(\theta_{2})-b(\theta_{1})a(\theta_{2}))c_{1}$ $=a(\theta)b(\theta_{2})-b(\theta)a(\theta_{2}),$ (27)
$(a(\theta_{2})b(\theta_{1})-b(\theta_{2})a(\theta_{1}))c_{2}$ $=a(\theta)b(\theta_{1})-b(\theta)a(\theta_{1}).$ (28)
If the expansion is done in such a non-orthogonal subspace, the minimization of $\langle\phi|U^{\dagger}(\vec{\theta})|H|U(\vec{\theta})|\phi\rangle$ transforms into a generalized eigenvalue problem (15) with matrices defined as
$\mathbf{H}=\left(\begin{array}[]{cc}\langle\phi|U^{\dagger}(\theta_{1})HU(\theta_{1})|\phi\rangle&\langle\phi|U^{\dagger}(\theta_{1})HU(\theta_{2})|\phi\rangle\\ \langle\phi|U^{\dagger}(\theta_{2})HU(\theta_{1})|\phi\rangle&\langle\phi|U^{\dagger}(\theta_{2})HU(\theta_{2})|\phi\rangle\end{array}\right).$ (31)
and
$\mathbf{S}=\left(\begin{array}[]{cc}1&\langle\phi|U^{\dagger}(\theta_{1})U(\theta_{2})|\phi\rangle\\ \langle\phi|U^{\dagger}(\theta_{2})U(\theta_{1})|\phi\rangle&1\end{array}\right).$ (34)
Solving this generalized eigenvalue problem allows us to determine the combination of $U(\theta_{1})|\phi\rangle$ and $U(\theta_{2})|\phi\rangle$ that corresponds to the lowest expectation value for $H$ in this two-configuration subspace.
Appendix C Energy gradient in ADAPT-GCIM framework
In the GCIM framework, the $k$-th eigenvalue of the projected $H$, $\epsilon_{k}$, can be represented as
$\epsilon_{k}=\frac{f^{T}_{k}\mathbf{H}f_{k}}{f^{T}_{k}\mathbf{S}f_{k}},$ (35)
where $f_{k}$ is the corresponding normalized eigenvectors, $f^{T}_{k}f_{k}=1$. Assume the working subspace comprises $k$ generating functions, $\{|\psi_{i}\rangle\}_{1\leq i\leq k}$, in which $k-1$ are generated by acting one Givens rotation on the reference, and only one is generated by acting the product form of Givens rotation on the reference, i.e.
$|\psi_{i}\rangle=\left\{\begin{array}[]{rc}G_{p_{i},q_{i}}(\theta_{i})|\phi_{0}\rangle,&1\leq i\leq k-1\\ \prod_{j=1}^{k-1}G_{p_{j},q_{j}}(\theta_{j})|\phi_{0}\rangle,&i=k\end{array}\right.\leavevmode\nobreak\ \leavevmode\nobreak\ .$ (38)
Then ${\bf H}$ and ${\bf S}$ are $k\times k$ matrices projected in this working subspace with their matrix elements being defined as
${\bf H}_{ij}$ $=\langle\psi_{i}|H|\psi_{j}\rangle,$ (39)
${\bf S}_{ij}$ $=\langle\psi_{i}|\psi_{j}\rangle,$ (40)
and the derivative of the eigenvalue $\epsilon_{k}$ with respect to the rotation $\theta_{s}$ can be computed through
$\frac{\partial\epsilon_{k}}{\partial\theta_{s}}=\frac{\langle\frac{\partial\mathbf{H}}{\partial\theta_{s}}\rangle_{k}\cdot\langle\mathbf{S}\rangle_{k}-\langle\mathbf{H}\rangle_{k}\cdot\langle\frac{\partial\mathbf{S}}{\partial\theta_{s}}\rangle_{k}}{\langle\mathbf{S}\rangle_{k}^{2}},$ (41)
with $\langle\mathbf{O}\rangle_{k}=f^{T}_{k}\mathbf{O}f_{k}$ for matrix $\mathbf{O}$. The matrix elements of $\frac{\partial\mathbf{S}}{\partial\theta_{s}}$ and $\frac{\partial\mathbf{H}}{\partial\theta_{s}}$ are the following,
  • if neither $|\psi_{i}\rangle$ nor $|\psi_{j}\rangle$ is parametrized by $\theta_{s}$
    $\frac{\partial(\mathbf{H}_{ij})}{\partial\theta_{s}}=\frac{\partial(\mathbf{S}_{ij})}{\partial\theta_{s}}=0,$ (42)
  • if only $|\psi_{i}\rangle$ is parametrized by $\theta_{s}$ and $i\neq j$
    $\frac{\partial(\mathbf{H}_{ij})}{\partial\theta_{s}}$ $=\langle\psi_{i}|(a_{q_{s}}^{\dagger}a_{p_{s}}-a_{p_{s}}^{\dagger}a_{q_{s}})H|\psi_{j}\rangle,$ (43)
    $\frac{\partial(\mathbf{S}_{ij})}{\partial\theta_{s}}$ $=\langle\psi_{i}|(a_{q_{s}}^{\dagger}a_{p_{s}}-a_{p_{s}}^{\dagger}a_{q_{s}})|\psi_{j}\rangle,$ (44)
  • if only $|\psi_{j}\rangle$ is parametrized by $\theta_{s}$ and $i\neq j$
    $\frac{\partial(\mathbf{H}_{ij})}{\partial\theta_{s}}$ $=\langle\psi_{i}|H(a_{p_{s}}^{\dagger}a_{q_{s}}-a_{q_{s}}^{\dagger}a_{p_{s}})|\psi_{j}\rangle,$ (45)
    $\frac{\partial(\mathbf{S}_{ij})}{\partial\theta_{s}}$ $=\langle\psi_{i}|(a_{p_{s}}^{\dagger}a_{q_{s}}-a_{q_{s}}^{\dagger}a_{p_{s}})|\psi_{j}\rangle,$ (46)
  • if both $|\psi_{j}\rangle$ are parametrized by $\theta_{s}$
    $\frac{\partial(\mathbf{H}_{ij})}{\partial\theta_{s}}$ $=\langle\psi_{i}|[H,(a_{p_{s}}^{\dagger}a_{q_{s}}-a_{q_{s}}^{\dagger}a_{p_{s}})]|\psi_{j}\rangle,$ (47)
    $\frac{\partial(\mathbf{S}_{ij})}{\partial\theta_{s}}$ $=0.$ (48)
As can be seen matrices $\frac{\partial\mathbf{S}}{\partial\theta_{s}}$ and $\frac{\partial\mathbf{H}}{\partial\theta_{s}}$ are very sparse with the non-zero elements only residing in the column and row whose index is equal to $p$, which can greatly reduce the computational complexity. The gradient employed in the proposed ADAPT-GCIM in Figure 3 can be considered as an approximation to the above gradient calculations, where only (47) and (48) are taken into consideration.
Appendix D ADAPT-VQE-GCIM and ADAPT-VQE-GCIM(1) algorithms
The main structure of ADAPT-VQE-GCIM in Alg. 1 is built upon a complete ADAPT-VQE process from Ref. 73. In fact, the essential use of ADAPT-VQE is to provide the unitary operators and phase values (except the phase for the latest selected ansätze) to form the bases for GCIM solvers. ADAPT-VQE-GCIM(1) only includes the selected single Givens rotations from all previous ADAPT iterations and the final product of the Givens rotations when ADAPT-VQE converges in the basis set.
Algorithm 1 ADAPT-VQE-GCIM
1:Operator pool $\{G_{l}\}_{(l)\in L}$ with $L$ the orbital index pool, initial state |$\ket{\psi_{(0)}^{VQE}}:=\ket{\phi_{HF}}$, VQE optimal parameter vector $\vec{\Theta}^{*}_{(0)}=\emptyset$, GCIM basis set $\mathbf{B}=\emptyset$, and Hamiltonian $H$
2:$k\leftarrow 0$
3:while convergence criteria is not met do
4:     $k\leftarrow k+1$
5:      Select operator: $G_{(k)}\leftarrow\max_{l\in L}\left\{\frac{\partial\epsilon}{\partial\theta_{l}}\big{|}_{\theta_{l}=0}=\bra{\psi^{VQE}_{(k-1)}}\left[H,A_{l}\right]\ket{\psi^{VQE}_{(k-1)}}\right\}$;
6:      Prepare VQE ansatz and optimize free parameters: |$\ket{\tilde{\psi}_{(k)}^{VQE}(\vec{\Theta}_{(k)})}=\prod_{i=1}^{k}G_{(i)}(\theta_{i})\ket{\phi_{HF}}$; $\vec{\Theta}^{*}_{(k)}=\operatorname*{arg\,min}_{\vec{\Theta}_{(k)}}\langle\tilde{\psi}_{(k)}^{VQE}(\vec{\Theta}_{(k)})|H\ket{\tilde{\psi}_{(k)}^{VQE}(\vec{\Theta}_{(k)})}$; |$\ket{\psi_{(k)}^{VQE}}\leftarrow\ket{\tilde{\psi}_{(k)}^{VQE}(\vec{\Theta}^{\ast}_{(k)})}$;
7:      Expand the basis set $B$: $\mathbf{B}\leftarrow G_{(k)}(\theta^{\ast}_{k})|\phi_{HF}\rangle$; $\mathbf{B}\leftarrow\ket{\psi_{(k)}^{VQE}}$;
8:      Compute matrices and solve for eigenvalues: $\mathbf{H}=\mathbf{B}^{T}H\mathbf{B}$; $\mathbf{S}=\mathbf{B}^{T}\mathbf{B}$; $\mathbf{H}f=\epsilon\mathbf{S}f$;
9:end while
10:return eigenvalues $\epsilon$ after convergence
Appendix E Implementation and resource estimation
For skew-Hermitian operators
$A_{p_{i},q_{i}}$ $=a_{p_{i}}^{\dagger}a_{q_{i}}-a_{q_{i}}^{\dagger}a_{p_{i}},\leavevmode\nobreak\ \leavevmode\nobreak\ q_{i}<p_{i}$ (49)
$A_{p_{i},r_{i},q_{i},s_{i}}$ $=a_{p_{i}}^{\dagger}a_{r_{i}}^{\dagger}a_{q_{i}}a_{s_{i}}-a_{s_{i}}^{\dagger}a_{q_{i}}^{\dagger}a_{r_{i}}a_{p_{i}},\leavevmode\nobreak\ \leavevmode\nobreak\ q_{i}<s_{i}<p_{i}<r_{i}$ (50)
their evolution in the qubit space can be expressed as
$\exp(\theta A_{p_{i},q_{i}})=\exp\bigg{(}-{\rm i}\frac{\theta}{2}(X_{q_{i}}Y_{p_{i}}-Y_{p_{i}}X_{q_{i}})\prod_{l=q_{i}+1}^{p_{i}-1}Z_{l}\bigg{)},$ (51)
$\exp(\theta A_{p_{i},r_{i},q_{i},s_{i}})=\exp\bigg{(}-{\rm i}\frac{\theta}{8}(X_{q_{i}}Y_{s_{i}}X_{p_{i}}X_{r_{i}}+Y_{q_{i}}X_{s_{i}}X_{p_{i}}X_{r_{i}}$
$\leavevmode\nobreak\ \leavevmode\nobreak\ +Y_{q_{i}}Y_{s_{i}}Y_{p_{i}}X_{r_{i}}+Y_{q_{i}}Y_{s_{i}}X_{p_{i}}Y_{r_{i}}-X_{q_{i}}X_{s_{i}}Y_{p_{i}}X_{r_{i}}-X_{q_{i}}X_{s_{i}}X_{p_{i}}Y_{r_{i}}$
$\leavevmode\nobreak\ \leavevmode\nobreak\ -Y_{q_{i}}X_{s_{i}}Y_{p_{i}}Y_{r_{i}}-X_{q_{i}}Y_{s_{i}}Y_{p_{i}}Y_{r_{i}})\prod_{k=p_{i}+1}^{r_{i}-1}Z_{k}\prod_{l=q_{i}+1}^{s_{i}-1}Z_{l}\bigg{)},$ (52)
under the Jordan-Widger mapping
$a_{p_{i}}=\frac{1}{2}(X_{p_{i}}+{\rm i}Y_{p_{i}})\prod_{k=0}^{p_{i}-1}Z_{k},\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ a_{p_{i}}^{\dagger}=\frac{1}{2}(X_{p_{i}}-{\rm i}Y_{p_{i}})\prod_{k=0}^{p_{i}-1}Z_{k}.$ (53)
A common way to implement unitary operators $\exp(\theta A_{p_{i},q_{i}})$ and $\exp(\theta A_{p_{i},r_{i},q_{i},s_{i}})$ in quantum circuits is to trotterize them into the product of the exponentials of single Pauli strings Nielsen and Chuang (2010). Without any circuit optimization, the standard method has the CNOT gate counts $4(p_{i}-q_{i})$ for Eq. (51) and $16(s_{i}-q_{i}+r_{i}-p_{i}+1)$ for Eq. (52Barkoutsos et al. (2018). In Ref. Yordanov et al. (2020), researchers reduce the CNOT gate counts to $12(p_{i}-q_{i})+$ 1 and $92(s_{i}-q_{i}+r_{i}-p_{i})+$ 9, respectively. To align with the quantum resource estimation for ADAPT-VQE in Ref. 63, we use the standard way to obtain the data in Figure 5. While the absolute numbers will drop significantly if we use the advanced implementations, the relative difference between ADAPT-VQE and ADAPT-GCIM still holds because both algorithms use exactly the same operator pool for each geometry in the simulation.
Alternatively, excitation operators can be directly implemented in terms of Givens rotation operators in circuits. For single excitations, Eq. (49), its corresponding Givens rotation operator for two adjacent indices is given in Figure 1, i.e., two CNOT gates. If the two indices are not adjacent, $p_{i}-q_{i}-1$ FSWAP gates will be employed on each side of the Givens rotation operator. If each FSWAP contains three CNOT gates, Hashim et al. (2021) the total number of CNOT gates for implementing a single excitation Givens rotation would be $46(p_{i}-q_{i})-$ 4. For double excitations, Eq. (50), a 4-qubit Givens rotation requires 14 CNOT gates assuming $p_{i},r_{i},q_{i},s_{i}$ are adjacent to each other Arrazola et al. (2022). Otherwise, $2(r_{i}-s_{i}+p_{i}-q_{i}-4)$ FSWAP ladders will be employed, which leads to $106(r_{i}-s_{i}+p_{i}-q_{i})-$ 10 CNOT gates for a double excitation Givens rotation. In Ref. 68, an efficient implementation for the general Givens rotation is discussed, where in the 4-qubit Hilbert space the Givens rotation applied to two adjacent spatial orbitals that preserves the number of $\alpha$ and $\beta$ spins, and the spin-squared operator $\hat{S}^{2}$ can be implemented using only 4 CNOT gates.
Appendix F Mitigation to the numerical instability
In the computation of $\mathbf{H}f=\epsilon\mathbf{S}f$ within GCIM framework, the overlap matrix $\mathbf{S}$ may be indefinite in experiments, especially when affected by a finite number of shots. To address this issue, one simple method is to add a small disturbance, such as $10^{-12}$, on all diagonal entries in $\mathbf{S}$. In our simulations, we adopted a more stable solution from Ref. 33. Since setting $\mathbf{S}_{ij}=\overline{\mathbf{S}}_{ji}$ ensures $\mathbf{S}$ is always Hermitian, the eigen-decomposition of $\mathbf{S}$ gives $\mathbf{S}=UDU^{\dagger}$ where $U$ is a unitary matrix composed of the eigenvectors of $\mathbf{S}$, and the $D$ is a diagonal matrix consisting of eigenvalues of $\mathbf{S}$. Not all diagonal entries of $D$ are positive in practice, so we only keep largest $L$ number of positive eigenvalues and truncate both $U$ and $D$ from $M\times M$ to $M\times L$ and $L\times L$, respectively. Then, the projection of $\mathbf{H}$ and $\mathbf{S}$ onto the subspace of the eigenspace of $\mathbf{S}$ is computed as following:
$\mathbf{H}_{trunc}$ $=U^{\dagger}_{trunc}\mathbf{H}U_{trunc}$ (54)
$\mathbf{S}_{trunc}$ $=U^{\dagger}_{trunc}\mathbf{S}U_{trunc}=D_{trunc}$ (55)
where $U_{trunc}$ and $D_{trunc}$ are the truncated matrices $U$ and $D$, respectively. Therefore, the reformulated generalized eigenvalue problem
$\mathbf{H}_{trunc}f_{trunc}=\epsilon_{trunc}D_{trunc}f_{trunc}$ (56)
replaces the original one.
To illustrate the effects of the quantum orthogonalization of the basis vectors, numerical simulations with classical orthogonalization on basis vectors in $LiH$ and both $H_{6}$ examples are conducted and demonstrated in Figure 6. It is noteworthy that there is no significant difference in the final accuracy. Combining this with the analysis in Appendix G, this subspace projection method could be effective enough to mitigate the potential numerical instability in near-future experiments.
Figure 6: Convergence performance before and after classical orthogonalization of the basis sets in ADAPT-VQE-GCIM and ADAPT-VQE for selected molecules. The “orth” in the labels mean the GCIM calculation was done with the “orthogonalized bases.”
Appendix G Randomness due to finite number of shots
Recall that in both ADAPT-GCIM and GCIM part of the ADAPT-VQE-GCIM, we compute project Hamiltonian, $\mathbf{H}$, and overlap matrix, $\mathbf{S}$, entry by entry
$\mathbf{H}_{ij}$ $=\langle\psi_{i}|H|\psi_{j}\rangle=\langle\phi_{HF}|B_{i}^{\dagger}HB_{j}|\phi_{HF}\rangle$ (57)
$\mathbf{S}_{ij}$ $=\langle\psi_{i}|\psi_{j}\rangle=\langle\phi_{HF}|B_{i}^{\dagger}B_{j}|\phi_{HF}\rangle,$ (58)
where $B_{i}$ and $B_{j}$ are short-hand notations for some products of Givens rotation matrices for simplicity. These expectations can be computed by converting $B_{i}^{\dagger}HB_{j}$ and $B_{i}^{\dagger}B_{j}$ to the sums of Pauli strings, respectively, then grouping the Pauli strings to do simultaneous measurements. The quantum and classical complexities and effects of shot noise have been explicitly discussed in Ref. 50.
Instead, in this section, we illustrate the shot noise calculation when employing Hadamard test for the expectation computation. Due to the randomness of the collapse of qubits, we will transfer $\mathbf{H}_{ij}$ to a random variable where the number of shots controls its variance.
Assume the Pauli decomposition of Hamiltonian $H=\sum_{k=1}^{N_{term}}c_{k}P_{k}$, $c_{k}\in\mathbb{R}$ for all $k$, has been provided, we have
$\mathbf{H}_{ij}=\sum_{k=1}^{N_{term}}c_{k}\langle\phi_{HF}|B_{i}^{\dagger}P_{k}B_{j}|\phi_{HF}\rangle.$ (59)
Without the loss of the generality, we only focus on $\mathbf{H}_{ij}$ since the calculation of $\mathbf{S}_{ij}$ can be regarded as a special case for (59) when $N_{term}=1$, $c_{k}=1$, and $P_{k}$ is an identity matrix. It is known that, Hadamard test gives
Pr(ancillary is |$\text{Pr}\left(\text{ancillary is }\ket{0}\right)$ $=\frac{1}{2}+\frac{1}{2}\text{Re}\langle\phi_{HF}|B_{i}^{\dagger}P_{k}B_{j}|\phi_{HF}\rangle$ (60)
Pr(ancillary is |$\text{Pr}\left(\text{ancillary is }\ket{1}\right)$ $=\frac{1}{2}-\frac{1}{2}\text{Re}\langle\phi_{HF}|B_{i}^{\dagger}P_{k}B_{j}|\phi_{HF}\rangle,$ (61)
after measuring the ancillary qubit in the circuit. Thus, a binary random variable can be defined by
$\Gamma^{(k)}_{l}=\begin{cases}1&\text{if ancillary is $\ket{0}$}\\ -1&\text{if ancillary is $\ket{1}$}\end{cases}$ (62)
in the $l^{th}$ shot when computing $\langle B_{i}^{\dagger}P_{k}B_{j}\rangle$ with
$p_{k}$ $:=\mathbb{E}\left[\Gamma^{(k)}_{l}\right]=\text{Re}\langle\phi_{HF}|B_{i}^{\dagger}P_{k}B_{j}|\phi_{HF}\rangle,$ (63)
$\text{Var}\left[\Gamma^{(k)}_{l}\right]$ $=\mathbb{E}\left[\left(\Gamma^{(k)}_{l}\right)^{2}\right]-\left(\mathbb{E}\left[\Gamma^{(k)}_{l}\right]\right)^{2}=1-p_{k}^{2},$ (64)
for all $l$ by assuming $\Gamma^{(k)}_{l}$’s are mutually independent and identically distributed. For simplicity, we omit the indices $i$ and $j$ for now. In other words, each measurement on the ancillary qubit in the Hadamard test circuit is modelled by sampling the random variable $\Gamma^{(k)}_{l}$.
We further construct another random variable, omitting indices $i$ and $j$ again,
$\Lambda^{(k)}=\sum_{l=1}^{N^{(k)}_{shot}}\Gamma^{(k)}_{l}$ (65)
where $N^{(k)}_{shot}$ is the number of shots for the Hadamard test circuit of $\langle B_{i}^{\dagger}P_{k}B_{j}\rangle$, and it is considered as a user-defined constant. Then, it follows that
$\frac{1}{2}\left(\Lambda^{(k)}+N^{(k)}_{shot}\right)\sim\text{Bin}\left(N^{(k)}_{shot},\frac{1}{2}+\frac{1}{2}p_{k}\right),$ (66)
by realizing $\frac{1}{2}+\frac{1}{2}\Gamma^{(k)}_{i}$ is a standard Bernoulli random variable. Further combining  (59) and (66), we obtain another discrete random variable
$\Xi^{(ij)}=\sum_{k=1}^{N_{term}}\frac{c_{k}}{N^{(k)}_{shot}}\Lambda^{(k)}=\sum_{k=1}^{N_{term}}\frac{c_{k}}{N^{(k)}_{shot}}\sum_{l=1}^{N^{(k)}_{shot}}\Gamma^{(k)}_{l}$ (67)
that satisfies
$\mathbb{E}\left[\Xi^{(ij)}\right]$ $=\sum_{k=1}^{N_{term}}\frac{c_{k}}{N^{(k)}_{shot}}N^{(k)}_{shot}p_{k}=\mathbf{H}_{ij}$ (68)
$\text{Var}\left[\Xi^{(ij)}\right]$ $=\sum_{k=1}^{N_{term}}\left(\frac{c_{k}}{N^{(k)}_{shot}}\right)^{2}N^{(k)}_{shot}\text{Var}\left[\Gamma^{(k)}_{l}\right]$
$=\sum_{k=1}^{N_{term}}\frac{c^{2}_{k}}{N^{(k)}_{shot}}\left(1-p_{k}^{2}\right).$ (69)
In this case, $\Xi^{(ij)}$ is an unbiased estimator of $\mathbf{H}_{ij}$ where the randomness is caused by the uncertainty from the collapse of the qubits during the measurement. The role of the number of shots, as (69) implies, is to govern the scale of the variance. The variance (69) goes to 0 when every $N_{shot}^{(k)}$ goes to infinity.
Chebyshev’s inequality shows that, for every $a>0$,
$\text{Pr}\left(\left|\Xi^{(ij)}-\mathbf{H}_{ij}\right|\geq a\right)\leq\frac{\text{Var}\left[\Xi^{(ij)}\right]}{a^{2}}=:\eta_{ij}.$ (70)
Setting $N^{(k)}_{shot}$ to $N_{shot}$ for all $i,j,k$, Table 6 provides the relation between $N_{shot}$ and $\max_{i,j}\,\eta_{ij}$ when $a$ is fixed at $10^{-4}$. However, the tightness of the bound is still unclear for this discrete variable, and this is not sufficient to illustrate the effects of the disturbance of the matrix entries on the eigenvalues. This ambiguity leads to numerical simulations.
Table 6: $N_{shot}$ for the various choices of $\max_{i,j}\,\eta_{ij}$ when $a=10^{-4}$.
Molecule $\max\,\eta_{ij}=5\%$ $\max\,\eta_{ij}=1\%$
$H4$ (square) $6.32\times 10^{10}$ $3.16\times 10^{11}$
$LiH$ $1.34\times 10^{11}$ $6.68\times 10^{11}$
[…]
Since it is very common to have huge $N^{(k)}_{shot}$ in practice, we can assume $\Lambda^{(k)}$ is closely approximated by
$\Lambda^{(k)}_{Gauss}\sim\text{N}\left(N^{(k)}_{shot}p_{k},\,\,N^{(k)}_{shot}\left(1-p_{k}^{2}\right)\right)$ (71)
and mutually independent for different $k$. Thus, $\Xi^{(ij)}$ is also approximated by a Gaussian random variable $\Xi^{(ij)}_{Gauss}$ with the same mean and variance as in (68) and (69), respectively.
The idea of importance sampling (IS) can be applied here to reduce the volatility without increasing the number of shots. By fixing the total number of shots for measuring all $P_{k}$’s as $N_{term}\cdot\tau$ with some large constant $\tau$, each individual Pauli string $P_{k}$ is allocated with
$N^{(k)}_{shot}:=\frac{|c_{k}|}{\sum_{k=1}^{N_{term}}|c_{k}|}(N_{term}\cdot\tau)$ (72)
number of shots. Under this setting, the random variable
$\Xi^{(ij)}_{IS}=\frac{\sum_{k=1}^{N_{term}}|c_{k}|}{N_{term}}\cdot\frac{1}{\tau}\sum_{k=1}^{N_{term}}\text{sign}(c_{k})\Lambda^{(k)}_{Gauss}$ (73)
is also Gaussian and unbiased with the mean $\mathbf{H}_{ij}$ and the variance
$\text{Var}\left[\Xi^{(ij)}_{IS}\right]=\frac{\sum_{k=1}^{N_{term}}|c_{k}|}{N_{term}}\cdot\frac{1}{\tau}\sum_{k=1}^{N_{term}}|c_{k}|\left(1-p_{k}^{2}\right).$ (74)
It is suggested to use larger $N^{k}_{shot}$ and $\tau$ for the overlap matrix $\mathbf{S}$ than $\mathbf{H}$ to improve the numerical stability. In the following simulations, $\mathbf{S}$ will always have 100 times larger $N^{k}_{shot}$ than $\mathbf{H}$.
The Monte Carlo simulations for squared $H_{4}$, $LiH$, and two $H_{6}$ geometries are conducted by treating each matrix entry as a normally distributed variable, following equations (68) and (69). We select a pair of $\mathbf{H}$ and $\mathbf{S}$ from a specific iteration of ADAPT-GCIM for each molecule, ensuring that the error in ground-state energy estimation is around $10^{-13}$ hartrees under the condition of an infinite number of shots. Although finite sampling gives challenges in reaching this level of accuracy, an accuracy of $10^{-4}$ to $10^{-5}$ hartrees is still attainable with the aid of importance sampling, as demonstrated in Figure 7.
Figure 7: Monte Carlo simulations for the effects of the shot noise on the accuracy of the selected iterations in ADAPT-GCIM on different molecules. The black and red-shaded areas are the $95\%$ confidence intervals estimated from 100 runs for each given $\tau$. When not using importance sampling, $N^{k}_{shot}=\tau$ for a projected Hamiltonian $\mathbf{H}$ and $N^{k}_{shot}=100\tau$ for an overlap matrix $\mathbf{S}$. With importance sampling, $N^{k}_{shot}$ follows (72) for $\mathbf{H}$ and 100 times larger for $\mathbf{S}$. Note that, given the same value of $\tau$, the total number of shots for estimating a matrix entry is the same with or without importance sampling..
Appendix H Intermittent and truncated optimization on ADAPT-GCIM
[…]
Figure 8: Convergence performance of ADAPT-GCIM with partially optimized parameters in computing the ground states. The label “ADAPT-GCIM(5,2)” means a parameter optimization is conducted for at most $2$ rounds in every $5$ ADAPT iterations.
Appendix I GCIM experiment on IBM quantum computer
[…]
$\mathbf{H}_{ij}=\bra{\psi_{i}}H\ket{\psi_{j}}=\bra{\phi_{HF}}U^{\dagger}_{i}(\theta)HU_{j}(\theta)\ket{\phi_{HF}},$
where $\theta=\pi/4$ as in other experiments, |$\ket{\phi_{HF}}$ is the HF state, we classically decompose the observable $U^{\dagger}_{j}(\pi/4)HU_{i}(\pi/4)$ into the linear combination of Pauli strings $\sum_{k}P_{k}$. So, quantum computer’s task was to evaluate a large number of $\bra{\phi_{HF}}P_{k}\ket{\phi_{HF}}$. By excluding the repeated Pauli strings across 465 entries, there are 8192 unique Pauli strings (8192 circuits) need to compute in this experiment. Of course, since $\bra{\phi_{HF}}P_{k}\ket{\phi_{HF}}=0$ for any $P_{k}$ that contains Pauli-$X$ or Pauli-$Y$, the quantum resources required can be further reduced by examining every elements in all 8192 Pauli strings. in exchange for classical running time.
We measured 1024 shots for each of 8192 circuits. With Qiskit Runtime’s default measurement error mitigation and zero-noise extrapolation (ZNE) Temme et al. (2017), the final error on ground-state energy estimation is $0.046$. Considering $\bra{\phi_{HF}}P\ket{\phi_{HF}}\in\{-1,0,1\}$ for any Pauli string $P$, we can efficiently build a filter to further mitigate quantum errors
$\bra{\phi_{HF}}P\ket{\phi_{HF}}=\begin{cases}1,&\text{if the noisy estimation $>0.2$}\\ -1&\text{if the noisy estimation $<-0.2$}\\ 0,&\text{otherwise}\end{cases}$ (75)
where the threshold $0.2$ is arbitrarily selected. Figure 9 clearly illustrate the distribution of 8192 expectations and the results after this filtering. With this problem-specific error mitigation method, error on ground-state energy estimation is reduced to $3.9\times 10^{-9}$ from the original $0.046$. In other words, we filtered out all the quantum errors and obtained the same result as in the classical simulation.
The purpose of the experiment is to examine the possibility of GCIM framework on the current quantum machines, although the heavy classical preprocessing undermines the potential quantum advantages. We will explore this possibility further with a more entangled circuit design based on Ref. 63; 67; 68; 69; 12; 71; 72; 82; 70; 83; 84.
Figure 9: Distribution of the expectations of 8192 unique Pauli strings, measured from ibm_osaka before and after the filtering in Eq. (75). Note that, raw data is already processed by the built-in measurement error mitigation and ZNE in Qiskit Runtime.